# Position Paper Are Normalizing Flows the Key to Unlocking the Exponential Mechanism? A Path through the Accuracy-Privacy Ceiling Constraining Differentially Private ML

### v3

@@ Appendix A Further Background on Differential Privacy
Differential privacy has many useful properties. First, post-processing is “free”, meaning if $f(\theta,X)$ is $(\varepsilon,\delta)$-DP, the same guarantee holds after composition with any function $h$; i.e., $h\circ f$ is also $(\varepsilon,\delta)$-DP (Dwork & Roth, 2014). Secondly, it satisfies composition theorems; the most basic of which says that if $f_{1}$ and $f_{2}$ satisfy $(\varepsilon_{1},\delta_{1})$-DP and $(\varepsilon_{2},\delta_{2})$-DP, respectively, then $f_{1}\circ f_{2}$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP (Dwork & Roth, 2014). The general problem of estimating the overall privacy bound (a single $\varepsilon,\delta)$) for a sequence of mechanisms $f_{i}(\theta_{i},X),i=1,...,m$ each furnished with a privacy guarantee $(\varepsilon_{i},\delta_{i})$-DP, is important for use.
As the basic composition theorem above holds, it is not a tight bound (better estimates may exist), and hence, a large body of work (Dwork, 2007, 2010; Kairouz et al., 2017; Abadi et al., 2016; Dong et al., 2021; Gopi et al., 2021) provide advanced composition theorems (tradeoffs of $\varepsilon$ and $\delta$ that depend on the $\varepsilon_{i},\delta_{i}$) that estimate this overall bound. Intuitively, as more information is released (as $m$ grows), the overall privacy guarantee gets worse ($\varepsilon$ and $\delta$ increase), and all advanced composition theorems entail an increase of $\delta$ for the better estimate of $\varepsilon$.
@@ Appendix B Further Background on Normalizing Flows
Following Berg et al. (Van Den Berg et al., 2018), a Sylvester Flow can be implemented by using the special case where $A=Q_{d\times d}R_{d\times m}$ and $B=\hat{R}_{m\times d}Q^{t}_{d\times d}$ with $Q$ orthogonal and $R,\hat{R}$ upper triangular. (This is a restriction in that it uses the same $Q$ for both $A$, $B$). To instantiate an orthogonal matrix $Q$, we define unit vectors $v$ and set $Q:=\prod_{v}I-2vv^{t}$ which is a product of Householder reflections. A well know fact is that Householder reflections are orthogonal, and any orthogonal matrix is as a product of Householder reflections (Sun & Bischof, 1995). Note this is obviated in the Planar Flow case since $A,B$ are simply $d$–dimensional vectors.
Appendix C Experiment Details
For DPSGD we use PRV accounting we use clipping parameter, max_grad_norm = 1 (Yousefpour et al., ; Yousefpour et al., 2021), $\delta=1\mathrm{e}{-5}$, roughly the inverse of the number of training data points, as are standard, and delta_error set to $\delta/1000$ (default). We compute the noise_multiplier parameter, which governs the variance of the Gaussian noise added to each gradient, to achieve the target privacy $(\varepsilon,\delta)$ based on the number of gradient steps (epochs and batch_size parameters) and the gradient clipping parameter max_grad_norm. The PRV Accounting method introduces an error term for $\varepsilon$ that bounds the error in the estimate of the privacy guarantee. Notably, eps_error =.01 is the default fidelity, and computations become too expensive or intractable for eps_error <.0001. Hence, given target epsilon $\bar{\varepsilon}$, we set eps_error to be as close as possible to $\bar{\varepsilon}/100$ within the range [.0001,.01].0001.01[.0001,.01][ .0001 , .01 ].
@@ Appendix C Experiment Details
While batch normalization is commonplace in non-private deep learning, it is not directly portable to the private setting as $z$-normalization learned in training must be shipped with the model (Yousefpour et al., 2021); hence, we do not use batch normalization for our private models, but we do provide results with vs. without batch normalization where applicable, namely for the non-private GRU-D model, Table 4. Details of the hyperparmeters used for each model can be found in our code base and vary per experiment and privacy level $\varepsilon$.
@@ C.1 MIMIC-III Details
Our experimental design follows a progression of research on MIMIC-III. Wang et al. (Wang et al., 2020) create benchmarks of needed healthcare predictive tasks from the MIMIC-III dataset, downloadable (from MIMIC-III v1.4, as all_hourly_data.h5) with the codebase (Wang et al., ). We use the binary prediction task benchmarks ICU Mortality and ICU Length of Stay ($>$ 3 days) from Wang et al.’s pipeline. These tasks differ only in target and share the feature set consisting of 104 measurement means per hour over a 24 hour period for 21,877 patients. Mortality is very imbalanced (07.4%/92.6%), while Length of Stay is balanced (47.1%/52.9%). We employ standardization ($z-$score normalization) for each feature, and it is applied independently per train/dev/test partition to prevent privacy issues.
@@
As an example use of the ICU Mortality and Length of Stay benchmarks, Wang et al. also present results of non-private LR and GRU-D (Che et al., 2018) models. Che et al. (Che et al., 2018) introduce the Gated Recurrent Unit with Decay (GRU-D) model, a deep neural network for time series learning tasks. GRU-D interpolates between the last seen value and the global mean so censored (missing) data does not need to be imputed in advance. Additionally, the hidden units of the GRU are decayed to zero with learned decay per component, which allows features that should have short influence (e.g., heart rate) vs. long influence (e.g., dosage/treatment) to have learned rate of diminishing value. Wang et al. provide non-private results for LR and GRU-D models on these two tasks as well as implementations.
@@
Suriyakumar et al. (Suriyakumar et al., 2021) replicate Wang et al.’s results on these two tasks for non-private baselines, then provide DPSGD (with RDP accountant) results as well. To the best of our knowledge, Suriyakumar et al. (Suriyakumar et al., 2021) present the state of the art in differentially private ML on these MIMIC-III benchmarks. Unfortunately, their code is not provided.
@@
Using the currently available data (MIMIC-III v1.4, as all_hourly_data.h5), and following preprocessing code and descriptions of Wang et al. we replicate these benchmarks as closely as possible. PRV accounting (Gopi et al., 2021) provides a better privacy bound than RDP, and we present comparative results across the two accounting methods. There are discrepancies between our dataset and Suriyakumar et al.’s description of the data, possibly because MIMIC-III was updated. The current MIMIC-III ICU Mortality and Length of Stay benchmarks have 89.1% missingness rate and 104 features per hour $\times$ 24 hours giving 2,496 columns. Suriyakumar et al. claim to have 78% missingness and only 78 features per hour, but they report the number of columns is reported as “24,69”, seemingly a typo (if the correct is 2,496, would entail 104 feature per hour). The number of rows matches exactly. The LR models entails privately learning these 2,497 parameters. We use RMSProp optimizer with a scheduler that reduces the learning rate for all LR models.
@@
For GRU-D, we follow Suriyakurmar et al. (Suriyakumar et al., 2021) and when needed the original GRU-D work of Che et al. (Che et al., 2018). Suriyakumar et al. (Suriyakumar et al., 2021) used batch normalization (BN) even when training privately with DPSGD, but this has since been revealed as a source of privacy leakage (Yousefpour et al., 2021). Hence, we do not use BN with our DPSDG implementations, but we do train a non-private model both with and without BN applied to the top regressor layer. We apply a dropout layer to the top regressor layer, treating the dropout probability as a hyperparameter. Unlike (Suriyakumar et al., 2021), we treat the hidden size as a hyperparameter for all experiments. The main motivation for this choice was that memory constraints on the GPU hardware used to perform ExpM+NF required that we used a hidden size no larger than 10. For the non-private and DPSGD implementations, we allow the hidden size to vary between 10 and 100 for our randomized grid search, so this constraint was only limiting for our new method. For training non-privately and with DPSGD, we use the Adam optimization method and for the NF model we use RMSProp with a scheduler. This is our the best possible replication of Wang et al. and Suriyakurmar et al.’s experiments save the choice to omit batch normalization.
@@
Notably, Che et al (Che et al., 2018) also provide non-private GRU-D results on MIMIC-III data for an ICU Mortality prediction task. Since Che et al. do not follow the preprocessing of Wang et al. and specifically claim to use 48 hours of data (while we, Wang, Suriyakurmar use 24 hours), we do not consider their results comparable.
@@
Appendix D Further Results
D.1 NF Distributions & Discussion
Recall our method is to train an NF model to approximate the ExpM density, which is a density over $\theta$, the parameters of our desired classifier. Our results use privacy values for ExpM+NF assuming the NF output is the ExpM distribution. While we cannot visualize the NF’s output distribution (because it is high dimensional), we can investigate the distribution of AUCs attained by ExpM+NF models for each $\varepsilon$ parameters. We present box plots in which indicate the distribution’s median (middle line), first (Q1) and third (Q3) quartiles (box), and extend 1.5 times the inner quartile range (Q1 - 1.5 IQR, Q3+ 1.5 IQR, whiskers). To see anomalous models arising from fortunate/unfortunate starting positions (different random seeds), we overlay each of the ten model’s means. We also include visualizations of the distribution of the AUC values produced by individual NFs. See Fig. 2. Results are similar across LR and GRU-D, and across both MIMIC-III prediction tasks. Randomness due to the model initialization is exhibited in the variance of each $\varepsilon$’s bar chart. Noise in the hyperparameter search process may influences the distribution trend across $\varepsilon$ (as a different hyperparmeter search is performed for each $\varepsilon$). For this method specifically, outlier NFs can produce bad results based on their randomized initializations. The increase in accuracy and privacy by ExpM+NF also affords a straightforward workaround, e.g., releasing ten models, each sampled from trained NFs with different starting seeds that are then used in a voting scheme.
Figure 2: For each $\varepsilon$ value, ten NF models are trained with identical hyperparmeters (but different starting seeds) and used to produce 1,000 samples each from the ExpM distribution. Right presents box plots of the 10,000 resulting classifiers’ AUC values. For each NF model’s 1,000 samples we compute the mean AUC, which are overlayed as dots. Left presents the AUC distribution from two NFs for four $\varepsilon$ values.
@@ D.1 NF Distributions & Discussion
D.2 Non-Private & DPSGD Advancements
Tables 3 & 4 present non-private LR & GRU-D results, respectively. We find show no significant difference in non-private or DPSGD trained models when using BCE vs. $\ell^{2}$ loss. (DPSGD BCE vs. $\ell^{2}$ results in Table 1). Note that Wang et al.’s LR used SciKit Learn (Pedregosa et al., 2011) while we use PyTorch.
Table 3: AUC scores for non-private LR model results for this and prior works on MIMIC-III tasks. Bold results are best per task. Underlined are the best result in this work.
(Wang et al., 2020) (Suriyakumar et al., 2021) Our $\ell^{2}$ loss Our BCE loss
@@ D.2 Non-Private & DPSGD Advancements
Table 4: AUC scores for non-private GRU-D model from this and prior works on MIMIC-III tasks. Bold results are best per task. Underlined results are best in this work per task. BN denotes batch normalization.
(Wang et al., 2020) (Suriyakumar et al., 2021) Our $\ell^{2}$ loss Our BCE loss
@@
Table 4 compares results of the GRU-D non-private baseline for previous works and ours, specifically including BCE vs $\ell^{2}$ loss and with batch normalization vs. without. We see nearly identical results in our results for each prediction task across the the two loss functions and with/without batch normalization, except for one instance, $\ell^{2}$ loss with batch normalization is worse. GRU-D non-private baseline results of our method compared to previous works reside in Table 4. Our results are very close but slightly below Wang et al. (Wang et al., 2020), yet better than Suriyakumar et al. (Suriyakumar et al., 2021). The gains in GRU-D’s performance of ours and Wang et al.’s results may be due to differences in data or other implementation details. Specifically, our hyperparameter search found that a hidden size of about 35-40 performed best, which is much smaller than the 67 hidden size of previous works (Suriyakumar et al., 2021; Che et al., 2018).
Our results show GRU-D, a much more expressive model than LR, does indeed surpass LR in AUC for non-private and each private training method in all experiments, respectively. This trend is seen in Che et al.(Che et al., 2018) and Wang et al. (Wang et al., 2020), but results are the inverse (LR is more accurate) in results of Suriyakumar et al. (Suriyakumar et al., 2021).
@@
Our replication of the same experiment produced LR models trained with DPSGD that greatly improve on this prior state-of-the-art. On ICU Mortality, Suriyakumar et al. (Suriyakumar et al., 2021) report 60% AUC for $(3.5,1\mathrm{e}{-5})$-DP, whereas we are able to achieve 76% AUC for the same $(3.5,1\mathrm{e}{-5})$-DP. Our DPSGD implementations are able to achieve $72-76\%$ AUC on this task for $0.5\leq\varepsilon\leq 10$ and maintain AUC above 60% until $\varepsilon<1\mathrm{e}{-1}$. For the Length of Stay task, Suriyakumar et al. (Suriyakumar et al., 2021) report 60% AUC for $(3.5,1\mathrm{e}{-5})$-DP, yet our models have 67% AUC for the same $\varepsilon$, which is only slightly below the non-private baseline. Again, we do not see an AUC below 60% until $\varepsilon<1\mathrm{e}{-1}$. Much of this improvement can be attributed to using the PRV Accounting method, which allowed us to use much less noise to achieve the same level of privacy.
@@
For the Length of Stay task, Suriyakumar et al. (Suriyakumar et al., 2021) report an AUC of 61% for GRU-D with $(3.5,1\mathrm{e}{-5})$-DP. We reach 70% AUC with DPSGD for the same level of privacy. This is better than the AUC reported by Suriyakumar et al. (Suriyakumar et al., 2021) for both the non-private LR and GRU-D models and we are able to maintain an AUC greater than 65% for $\varepsilon>1\mathrm{e}{-2}$ with DPSGD. As we discussed in the prior section, much of the gains in AUC for GRU-D with DPSGD are likely due using the PRV Accounting method. Some of the gains could also be attributed to the same causes that provided the large gains we made for the non-private GRU-D model. Suriyakumar et al. (Suriyakumar et al., 2021) show a sharp decrease in AUC when training GRU-D with DPSGD on ICU Mortality task, even in the “low privacy” case with $\varepsilon=3\mathrm{e}5$. They report 59% AUC for this high $\varepsilon$ and 53% AUC for the “strong privacy” case, $(3.5,1\mathrm{e}{-5})$-DP, which is not much better than random guessing. Our DPSGD PRV implementation of GRU-D achieves median 82.7% AUC for $(3.5,1\mathrm{e}{-5})$-DP using GRU-D on the Mortality prediction task. This is comparable to the best non-private LR model that we achieved, and is better than the performance (Suriyakumar et al., 2021) reported for non-private GRU-D. Our DPSGD results achieve AUC greater than 60% for each $\varepsilon>1\mathrm{e}{-2}$ in this prediction task.
D.3 Training Time Details
The details of this subsection correspond with performance results in Sec. 4.2. As all our models are implemented in PyTorch, we use PyTorch’s benchmark suite over ten runs on a single NVIDIA A100 80GB GPU and 128 CPUs with 2TB of shared RAM. For DPSGD and ExpM+NF, we additionally take the mean across the various privacy levels, noting that some $\varepsilon$ may have hyperparameters that led it to take more or less time than this reported mean. For the non-private GRU-D implementation, we report mean training time with batch normalization which took less time. In the case of the non-private BCE loss, the number of epochs and batch size were larger, likely attributing to the increased training time.
Table 5: Length of Stay Benchmark Timing Results

### v4

@@ Appendix A Further Background on Differential Privacy
Differential privacy has many useful properties. First, post-processing is “free”, meaning if $f(\theta,X)$ is $(\varepsilon,\delta)$-DP, the same guarantee holds after composition with any function $h$; i.e., $h\circ f$ is also $(\varepsilon,\delta)$-DP [11]. Secondly, it satisfies composition theorems; the most basic of which says that if $f_{1}$ and $f_{2}$ satisfy $(\varepsilon_{1},\delta_{1})$-DP and $(\varepsilon_{2},\delta_{2})$-DP, respectively, then $f_{1}\circ f_{2}$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP [11]. The general problem of estimating the overall privacy bound (a single $\varepsilon,\delta)$) for a sequence of mechanisms $f_{i}(\theta_{i},X),i=1,...,m$ each furnished with a privacy guarantee $(\varepsilon_{i},\delta_{i})$-DP, is important for use.
As the basic composition theorem above holds, it is not a tight bound (better estimates may exist), and hence, a large body of work [9, 10, 19, 1, 8, 16] provide advanced composition theorems (tradeoffs of $\varepsilon$ and $\delta$ that depend on the $\varepsilon_{i},\delta_{i}$) that estimate this overall bound. Intuitively, as more information is released (as $m$ grows), the overall privacy guarantee gets worse ($\varepsilon$ and $\delta$ increase), and all advanced composition theorems entail an increase of $\delta$ for the better estimate of $\varepsilon$.
@@ Appendix B Further Background on Normalizing Flows
Following Berg et al. [37], a Sylvester Flow can be implemented by using the special case where $A=Q_{d\times d}R_{d\times m}$ and $B=\hat{R}_{m\times d}Q^{t}_{d\times d}$ with $Q$ orthogonal and $R,\hat{R}$ upper triangular. (This is a restriction in that it uses the same $Q$ for both $A$, $B$). To instantiate an orthogonal matrix $Q$, we define unit vectors $v$ and set $Q:=\prod_{v}I-2vv^{t}$ which is a product of Householder reflections. A well know fact is that Householder reflections are orthogonal, and any orthogonal matrix is as a product of Householder reflections [35]. Note this is obviated in the Planar Flow case since $A,B$ are simply $d$–dimensional vectors.
Appendix C Further Details of ExpM+NF
Our utility must be $u=\phi\circ L$, a decreasing function of the loss. While any strictly decreasing $\phi$ theoretically ensures identical optimization, in practice $\phi$ must separate the near-optimal $\theta$ from those that are far from optimal. This means the derivative must be a sufficiently large negative value over an appropriate range of $L(\theta)$ to discriminate high versus low utility $\theta$. In particular, using a negative sigmoid function, which is flat ($\phi^{\prime}\approxeq 0$) except in a limited neighborhood, is unwise, unless we know which range of $L(\theta)$ will contain the optimal values a priori and shift $\phi$ to discriminate these values. Lastly, we note that choice of $\phi$ effects the sensitivity, and therefore may be designed to assist in producing the required sensitivity bound.
When $s$ is large, it diminishes influence of $u$ on $p_{ExpM}$ and increasing the variance. This leads to higher likelihood of sampling less optimal $\theta$. Since $u,s$ are intimately tied to the loss function $L,$ bounding $s$ is generally required per loss function and can depend on the model outputs. Hence, if for a particular model and loss function a strong bound can be made, this can give ExpM+NF a boost, and conversely, if only a weak or no sensitivity bound can be identified, ExpM+NF will be hindered or useless.
Appendix D MIMIC-III ICU Benchmarks Experiment Details
For DPSGD we use PRV accounting we use clipping parameter, max_grad_norm = 1 [42, 43], $\delta=1\mathrm{e}{-5}$, roughly the inverse of the number of training data points, as are standard, and delta_error set to $\delta/1000$ (default). We compute the noise_multiplier parameter, which governs the variance of the Gaussian noise added to each gradient, to achieve the target privacy $(\varepsilon,\delta)$ based on the number of gradient steps (epochs and batch_size parameters) and the gradient clipping parameter max_grad_norm. The PRV Accounting method introduces an error term for $\varepsilon$ that bounds the error in the estimate of the privacy guarantee. Notably, eps_error =.01 is the default fidelity, and computations become too expensive or intractable for eps_error <.0001. Hence, given target epsilon $\bar{\varepsilon}$, we set eps_error to be as close as possible to $\bar{\varepsilon}/100$ within the range [.0001,.01].0001.01[.0001,.01][ .0001 , .01 ].
@@ Appendix D MIMIC-III ICU Benchmarks Experiment Details
While batch normalization is commonplace in non-private deep learning, it is not directly portable to the private setting as $z$-normalization learned in training must be shipped with the model [43]; hence, we do not use batch normalization for our private models, but we do provide results with vs. without batch normalization where applicable, namely for the non-private GRU-D model, Table 4. Details of the hyperparmeters used for each model can be found in our code base and vary per experiment and privacy level $\varepsilon$.
@@ D.1. MIMIC-III Dataset Details
MIMIC-III [18] includes records for a diverse group of critical care patients at a medical center from 2001–2012. Our experimental design follows a progression of research on MIMIC-III. Wang et al. [40] create benchmarks of needed healthcare predictive tasks from the MIMIC-III dataset, downloadable (from MIMIC-III v1.4, as all_hourly_data.h5) with the codebase [39]. We use the binary prediction task benchmarks ICU Mortality and ICU Length of Stay ($>$ 3 days) from Wang et al.’s pipeline. These tasks differ only in target and share the feature set consisting of 104 measurement means per hour over a 24 hour period for 21,877 patients. Mortality is very imbalanced (07.4%/92.6%), while Length of Stay is balanced (47.1%/52.9%). We employ standardization ($z-$score normalization) for each feature, and it is applied independently per train/dev/test partition to prevent privacy issues.
@@
As an example use of the ICU Mortality and Length of Stay benchmarks, Wang et al. also present results of non-private LR and GRU-D [4] models. Che et al. [4] introduce the Gated Recurrent Unit with Decay (GRU-D) model, a deep neural network for time series learning tasks. GRU-D interpolates between the last seen value and the global mean so censored (missing) data does not need to be imputed in advance. Additionally, the hidden units of the GRU are decayed to zero with learned decay per component, which allows features that should have short influence (e.g., heart rate) vs. long influence (e.g., dosage/treatment) to have learned rate of diminishing value. Wang et al. provide non-private results for LR and GRU-D models on these two tasks as well as implementations.
@@
Suriyakumar et al. [36] replicate Wang et al.’s results on these two tasks for non-private baselines, then provide DPSGD (with RDP accountant) results as well. To the best of our knowledge, Suriyakumar et al. [36] present the state of the art in differentially private ML on these MIMIC-III benchmarks. Unfortunately, their code is not provided.
@@
Using the currently available data (MIMIC-III v1.4, as all_hourly_data.h5), and following preprocessing code and descriptions of Wang et al. we replicate these benchmarks as closely as possible. PRV accounting [16] provides a better privacy bound than RDP, and we present comparative results across the two accounting methods. There are discrepancies between our dataset and Suriyakumar et al.’s description of the data, possibly because MIMIC-III was updated. The current MIMIC-III ICU Mortality and Length of Stay benchmarks have 89.1% missingness rate and 104 features per hour $\times$ 24 hours giving 2,496 columns. Suriyakumar et al. claim to have 78% missingness and only 78 features per hour, but they report the number of columns is reported as “24,69”, seemingly a typo (if the correct is 2,496, would entail 104 feature per hour). The number of rows matches exactly. The LR models entails privately learning these 2,497 parameters. We use RMSProp optimizer with a scheduler that reduces the learning rate for all LR models.
@@
For GRU-D, we follow Suriyakurmar et al. [36] and when needed the original GRU-D work of Che et al. [4]. Suriyakumar et al. [36] used batch normalization (BN) even when training privately with DPSGD, but this has since been revealed as a source of privacy leakage [43]. Hence, we do not use BN with our DPSDG implementations, but we do train a non-private model both with and without BN applied to the top regressor layer. We apply a dropout layer to the top regressor layer, treating the dropout probability as a hyperparameter. Unlike [36], we treat the hidden size as a hyperparameter for all experiments. The main motivation for this choice was that memory constraints on the GPU hardware used to perform ExpM+NF required that we used a hidden size no larger than 10. For the non-private and DPSGD implementations, we allow the hidden size to vary between 10 and 100 for our randomized grid search, so this constraint was only limiting for our new method. For training non-privately and with DPSGD, we use the Adam optimization method and for the NF model we use RMSProp with a scheduler. This is our the best possible replication of Wang et al. and Suriyakurmar et al.’s experiments save the choice to omit batch normalization.
@@
Notably, Che et al [4] also provide non-private GRU-D results on MIMIC-III data for an ICU Mortality prediction task. Since Che et al. do not follow the preprocessing of Wang et al. and specifically claim to use 48 hours of data (while we, Wang, Suriyakurmar use 24 hours), we do not consider their results comparable.
@@
Appendix E Further MIMIC-III Experiment Results
E.1. NF Distributions & Discussion
Recall our method is to train an NF model to approximate the ExpM density, which is a density over $\theta$, the parameters of our desired classifier. Our results use privacy values for ExpM+NF assuming the NF output is the ExpM distribution. While we cannot visualize the NF’s output distribution (because it is high dimensional), we can investigate the distribution of AUCs attained by ExpM+NF models for each $\varepsilon$ parameters. We present box plots which indicate the distribution’s median (middle line), first (Q1) and third (Q3) quartiles (box), and extend 1.5 times the inner quartile range (Q1 - 1.5 IQR, Q3+ 1.5 IQR, whiskers). To see anomalous models arising from fortunate/unfortunate starting positions (different random seeds), we overlay each of the ten model’s means. We also include visualizations of the distribution of the AUC values produced by individual NFs. See Fig. 7. Results are similar across LR and GRU-D, and across both MIMIC-III prediction tasks. Randomness due to the model initialization is exhibited in the variance of each $\varepsilon$’s bar chart. Noise in the hyperparameter search process may influences the distribution trend across $\varepsilon$ (as a different hyperparmeter search is performed for each $\varepsilon$). For this method specifically, outlier NFs can produce bad results based on their randomized initializations. The increase in accuracy and privacy by ExpM+NF also affords a straightforward workaround, e.g., releasing ten models, each sampled from trained NFs with different starting seeds that are then used in a voting scheme.
Figure 7. For GRU-D model on ICU Mortality task, for each $\varepsilon$ value, ten NF models are trained with identical hyperparmeters (but different starting seeds) and used to produce 1,000 samples each from the ExpM distribution. Left presents box plots of the 10,000 resulting classifiers’ AUC values. For each NF model’s 1,000 samples we compute the mean AUC, which are overlayed as dots. Right presents the AUC distribution from two NFs for four $\varepsilon$ values.
@@ E.1. NF Distributions & Discussion
E.2. Non-Private & DPSGD Advancements
Tables 3 & 4 present non-private LR & GRU-D results, respectively. We find show no significant difference between BCE and $\ell^{2}$ loss for non-private and DPSGD-trained models. (DPSGD BCE vs. $\ell^{2}$ results in Table 1). Note that Wang et al.’s LR used SciKit Learn [30] while we use PyTorch.
Table 3. AUC scores for non-private LR model results for this and prior works on MIMIC-III tasks. Bold results are best per task. Underlined are the best result in this work.
[40] [36] Our $\ell^{2}$ loss Our BCE loss
@@ E.2. Non-Private & DPSGD Advancements
Table 4. AUC scores for non-private GRU-D model from this and prior works on MIMIC-III tasks. Bold results are best per task. Underlined results are best in this work per task. BN denotes batch normalization.
[40] [36] Our $\ell^{2}$ loss Our BCE loss
@@
Table 4 compares results of the GRU-D non-private baseline for previous works and ours, specifically including BCE vs $\ell^{2}$ loss and with batch normalization vs. without. We see nearly identical results in our results for each prediction task across the the two loss functions and with/without batch normalization, except for one instance, $\ell^{2}$ loss with batch normalization is worse. GRU-D non-private baseline results of our method compared to previous works reside in Table 4. Our results are very close but slightly below Wang et al. [40], yet better than Suriyakumar et al. [36]. The gains in GRU-D’s performance of ours and Wang et al.’s results may be due to differences in data or other implementation details. Specifically, our hyperparameter search found that a hidden size of about 35-40 performed best, which is much smaller than the 67 hidden size of previous works [36, 4].
Our results show GRU-D, a much more expressive model than LR, does indeed surpass LR in AUC for non-private and each private training method in all experiments, respectively. This trend is seen in Che et al.[4] and Wang et al. [40], but results are the inverse (LR is more accurate) in results of Suriyakumar et al. [36].
@@
Our replication of the same experiment produced LR models trained with DPSGD that greatly improve on this prior state-of-the-art. On ICU Mortality, Suriyakumar et al. [36] report 60% AUC for $(3.5,1\mathrm{e}{-5})$-DP, whereas we are able to achieve 76% AUC for the same $(3.5,1\mathrm{e}{-5})$-DP. Our DPSGD implementations are able to achieve $72-76\%$ AUC on this task for $0.5\leq\varepsilon\leq 10$ and maintain AUC above 60% until $\varepsilon<1\mathrm{e}{-1}$. For the Length of Stay task, Suriyakumar et al. [36] report 60% AUC for $(3.5,1\mathrm{e}{-5})$-DP, yet our models have 67% AUC for the same $\varepsilon$, which is only slightly below the non-private baseline. Again, we do not see an AUC below 60% until $\varepsilon<1\mathrm{e}{-1}$. Much of this improvement can be attributed to using the PRV Accounting method, which allowed us to use much less noise to achieve the same level of privacy.
@@
For the Length of Stay task, Suriyakumar et al. [36] report an AUC of 61% for GRU-D with $(3.5,1\mathrm{e}{-5})$-DP. We reach 70% AUC with DPSGD for the same level of privacy. This is better than the AUC reported by Suriyakumar et al. [36] for both the non-private LR and GRU-D models and we are able to maintain an AUC greater than 65% for $\varepsilon>1\mathrm{e}{-2}$ with DPSGD. As we discussed in the prior section, much of the gains in AUC for GRU-D with DPSGD are likely due using the PRV Accounting method. Some of the gains could also be attributed to the same causes that provided the large gains we made for the non-private GRU-D model. Suriyakumar et al. [36] show a sharp decrease in AUC when training GRU-D with DPSGD on ICU Mortality task, even in the “low privacy” case with $\varepsilon=3\mathrm{e}5$. They report 59% AUC for this high $\varepsilon$ and 53% AUC for the “strong privacy” case, $(3.5,1\mathrm{e}{-5})$-DP, which is not much better than random guessing. Our DPSGD PRV implementation of GRU-D achieves median 82.7% AUC for $(3.5,1\mathrm{e}{-5})$-DP using GRU-D on the Mortality prediction task. This is comparable to the best non-private LR model that we achieved, and is better than the performance [36] reported for non-private GRU-D. Our DPSGD results achieve AUC greater than 60% for each $\varepsilon>1\mathrm{e}{-2}$ in this prediction task.
E.3. Training Time Details
The details of this subsection correspond with performance results in Sec. 4.1. Note that either by choice or a bug, PyTorch does not expose the ten runs’ times, nor statistics beside their mean https://github.com/pytorch/pytorch/issues/106801. In the case of the non-private BCE loss, the number of epochs and batch size were larger, likely attributing to the increased training time. Raw results for each training method, task, and $\varepsilon$ appear in the repository.
Table 5. Bayesian Inference and MAP Estimate Results using Logistic Regression on the Mortality ICU Benchmark
$\tilde{\varepsilon}$ 0.001 0.01 0.1 1
Bayesian AUC 0.79 0.81 0.80 0.80
MAP AUC 0.79 0.81 0.78 0.80
Table 6. Length of Stay Benchmark Timing Results